9:20 a.m. Morning Update

Per Texas Brine, drilling continues to remain on schedule. This weekend, activity at the drilling site will include dismantling and moving out the current (on-site) rig and replacing it with a snubbing rig which will continue the final drilling operation into the cavern. You may see heavier traffic as you did when the current rig was moved onto the site. We do not anticipate a complete closure of Highway 70; however, there may be intermittent (and very short) traffic stops to move equipment in and out.

DNR contractors will continue with the geo-probes on the north side of LA-70. Also, a sonic rig will be arriving this afternoon to drill into the water aquifer on the south side of LA-70 on Dugas & LeBlanc property. The operation is expected to be completed sometime tomorrow and information of findings should be known shortly thereafter and posted as soon as possible.